OpenStudy (anonymous):

solve for r r^2+3r=8 Example r^2+3r=1 r=-3+√13/2,-3-√13/2

OpenStudy (anonymous):

looks like the quadratic equation might help here

OpenStudy (anonymous):

r^2 +3r -8 =0 a=1 b=3 c=-8 Use the quadratic formula to get your answer(s) for x.

OpenStudy (anonymous):

The quadratic equation is where my problem lies

OpenStudy (anonymous):

Do you know what the quadratic formula is?

OpenStudy (anonymous):

((-1) +/- sqrt ((3)^2-(4)(1)(-8)))/((2)(1))

OpenStudy (anonymous):

(-1 +/- sqrt (41))/2 ***THAT MEANS*** x = (-1 + sqrt(41))/2 x = (-1 - sqrt(41))/2

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

This is what you do whenever factoring is impossible.

OpenStudy (anonymous):

Or if you're too lazy to factor :-)... like I usually am LOL

OpenStudy (anonymous):

so how would i put x^2+7x=10 into the quadratic equation

OpenStudy (anonymous):

Okay so your format is going to be ax^2+bx+c So with x^2+7x=10, you always always always get everything on the left side. You're setting it equal to zero. x^2+7x-10=0 Now what are a, b, and c?

OpenStudy (anonymous):

sorry typed problem wrong x^2+7x-10=0

OpenStudy (anonymous):

k... so what do you think a, b, and c are?

OpenStudy (anonymous):

a is 1 b is 7 c is 10

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

I'll ask you one now. -x^2+2x+3 What is "a" in this case?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

my problem is keeping it all in order

OpenStudy (anonymous):

Now if there's a term, say b=-2 and you need to plug it into your formula, the first thing you're putting in is -b. So how to do that: -(-2). So that will be a +2.

OpenStudy (anonymous):

I understand. And I would recommend to do the problem in as many steps as you need to. And to slow down and pay close attention. I took an engineering problems course last quarter and I made a lot of mistakes when I didn't pay attention :-)

OpenStudy (anonymous):

You can try that problem and we'll compare answers if you want.

OpenStudy (anonymous):

ok is it -7+√89/2,-7-√89/2

OpenStudy (anonymous):

That's what I got too. You know what you're doing!

OpenStudy (anonymous):

You have no idea how much you have helped me thanks again!!!!!

OpenStudy (anonymous):

I'm glad :-) So step 1: Get everything on the left side so that it's set equal to zero. step 2: write down what your a, b, and c values are step 3: plug in to the formula and go slow

OpenStudy (anonymous):

I truly cannot thank you enough I have been struggling for days

OpenStudy (anonymous):

I would recommend Paul's online notes as an alternative to your book. Most algebra books truly suck. http://tutorial.math.lamar.edu/Classes/Alg/SolveQuadraticEqnsII.aspx